Given 2+x/5= x/7. What is the first step to solve this proportion?
Answer by Alan3354(69443)

Given 2+x/5= x/7. What is the first step to solve this proportion?
--------------
Eliminate fractions by multiplying thru by 35.
The "shortcut" is to cross multiply:
(2+x)*7 = 5x